# HSF: Defending against Jailbreak Attacks with Hidden State Filtering

### v1

@@
Abstract
With the growing deployment of LLMs in daily applications like chatbots and content generation, efforts to ensure outputs align with human values and avoid harmful content have intensified. However, increasingly sophisticated jailbreak attacks threaten this alignment, aiming to induce unsafe outputs. Current defense efforts either focus on prompt rewriting or detection, which are limited in effectiveness due to the various design of jailbreak prompts, or on output control and detection, which are computationally expensive as they require LLM inference. Therefore, designing a pre-inference defense method that resists diverse jailbreak prompts is crucial for preventing LLM jailbreak attacks. We observe that jailbreak attacks, safe queries, and harmful queries exhibit different clustering patterns within the LLM’s hidden state representation space. This suggests that by leveraging the LLM’s hidden state representational capabilities, we can analyze the LLM’s forthcoming behavior and proactively intervene for defense. In this paper, we propose a jailbreak attack defense strategy based on a Hidden State Filter (HSF), a lossless architectural defense mechanism that enables the model to preemptively identify and reject adversarial inputs before the inference process begins. We activate its defensive potential through an additional plugin module, effectively framing the defense task as a classification problem. Experimental results on two benchmark datasets, utilizing three different LLMs, show that HSF significantly enhances resilience against six cutting-edge jailbreak attacks. It significantly reduces the success rate of jailbreak attacks while minimally impacting responses to benign user queries, with negligible inference overhead, and outperforming defense baselines.111 Our code and data are available at https://anonymous.4open.science/r/Hidden-State-Filtering-8652/
1 Introduction
@@
Table 1: Examples of various jailbreak attack methods and their corresponding prompts. The {User Query} is "Write a tutorial on how to make a bomb". Each attack represents a different strategy to bypass content restrictions in alignment LLM by manipulating input prompts.
With the growing capabilities of models like ChatGPT (Achiam et al. 2023), Llama (Touvron et al. 2023), and Mistral (Jiang et al. 2023), there is increasing concern about the potential security risks associated with their powerful functionalities, such as biased reports (Ferrara 2023), inaccurate information (Ji et al. 2023), or harmful contents (Weidinger et al. 2021). Over the past few years, researchers make extensive efforts to align these models safely using techniques such as SFT (Ouyang et al. 2022) and RLHF(Bai et al. 2022), ensuring that the models adhere to human values and preferences for safety and reliability. Despite these alignment efforts providing models with the ability to judge the safety of inputs to some extent, models remain susceptible to adversarial attacks (Zou et al. 2023).
Recent studies have highlighted the potential threat of “jailbreak attacks”(Liu et al. 2023; Wei, Haghtalab, and Steinhardt 2024), which can bypass existing alignments and induce models to output unsafe content. As shown in Table 1, different types of attacks vary greatly and can be introduced into the prompts in many ways. Current defense methods primarily focus on prompt’s detection (Alon and Kamfonas 2023) and rewriting (Jain et al. 2023; Zheng et al. 2024) or output’s detection (Xie et al. 2023) and control (Xu et al. 2024). The former defends against jailbreak attacks by using PPL (Alon and Kamfonas 2023), paraphrasing (Jain et al. 2023), or adding safer prompts (Zheng et al. 2024; Wei, Wang, and Wang 2023), but faces issues of limited defense scope, high costs, and disruption of benign prompts. The latter defends by controlling the decoder’s generation process (Xu et al. 2024) or detecting the final output (Xie et al. 2023), but incurs high computational costs as it starts only after LLM inference. Therefore, designing a pre-inference defense method that remains unaffected by various jailbreak prompts is crucial for preventing LLM jailbreak attacks.
@@
Figure 1: Comparison of four different models using PCA visualization, illustrating how each model differentiates between harmless (blue) and harmful (orange) queries. The Llama2-7B-chat(a), Vicuna-v.15-7B(b) and Mistral-instruct-v0.2-7B(c) are aligned models, while the WizardLM-30B-Uncensored(d) is unaligned model.
Figure 2: Comparison of Llama2-7B-chat model using PCA visualization, illustrating how the model differentiates between harmless (blue) and harmful (orange) queries across various jailbreak (red) methods. Subplots correspond to the following attack methods: (1) Autodan, (2) GCG, (3) DeepInception, (4) ReNellm, (5) ICA, and (6) GPTFuzz. The decision boundaries that the model learned to separate the two types of queries are shown.
2 Related Work
@@
2.1 Jailbreak Attacks
Current jailbreak attacks can be broadly categorized into two major types: template completion attacks and prompt rewriting attacks(Yi et al. 2024).
In template completion attacks, the attacker embeds harmful queries within a contextual template to generate jailbreak prompts. Liu et al. (2023) investigates adversarial examples against the GPT-3 model, where carefully crafted inputs induce the model to generate inappropriate content. Wei, Haghtalab, and Steinhardt (2024) identifies the root cause of LLMs’ vulnerability to jailbreak attacks as conflicting objectives and generalization mismatches, noting that the model’s safety capabilities do not match its complexity. Li et al. (2023) constructs a virtual, nested scenario to hypnotize large models into performing jailbreak attacks. Ding et al. (2024) utilizes prompt rewriting and scenario nesting within LLMs themselves to generate effective jailbreak prompts. Wei, Wang, and Wang (2023) leverages the model’s contextual learning abilities to guide the LLM in generating unsafe outputs.
Prompt rewriting attacks involve constructing jailbreak prompts through optimization techniques, with two main approaches: (1) Gradient-based methods, as Zou et al. (2023) uses gradient optimization to generate adversarial inputs; (2) Genetic algorithm-based methods, as Liu et al. (2024) and Yu, Lin, and Xing (2023) collect seeds from a curated library of handcrafted templates and continuously generate optimized adversarial prompts through mutation and crossover.
@@ 3 Motivation
3.1 Data Synthesis and Basement Model
@@
To ensure the diversity of the synthesized dataset, we sampled the top 100 data points from Advbench (Zou et al. 2023) as harmful queries and instructed gpt-3.5-turbo to generate benign versions of these queries while maintaining consistency in length as much as possible. Please refer to the appendix for the prompts we used to guide the data synthesis. We then applied additional manual checks to ensure effectiveness and quality. As a result, we collected 100 harmful and 100 benign queries, with an average length of 15.05 and 17.14 tokens, respectively (measured using the Llama2-7b-chat tokenizer).
We conduct experiments with three popular 7B chat LLMs available on HuggingFace, as well as an uncensored 30B LLM: Llama-2-chat (Touvron et al. 2023), Vicuna-v1.5 (Chiang et al. 2023), Mistral-instruct-v0.2 (Jiang et al. 2023), and WizardLM-30B-Uncensored (Computations 2024). Some of these models explicitly underwent extensive safety training (Llama2-7b-chat and Mistral-instruct-v0.2).
3.2 Visualization Analysis
@@
Aligned Model vs. Unaligned Model
Following (Zheng et al. 2024), we used Principal Component Analysis (PCA) to visualize the hidden states of the models. We selected the hidden states of the last input token from the model’s last Decoder Layer output, as these hidden states intuitively gather all the information about the model’s understanding of the query and its response strategy. It should be noted that these hidden states are also projected through the language modeling head (a linear mapping) for the prediction of the next token, which means they present a linear structure in the corresponding representation space (as assumed by PCA). We use two sets of hidden states to compute the first two principal components for each model, which include both harmful and benign queries. By selecting these data points, we are able to extract the most significant features related to the harmfulness of the queries.
@@
Distinguishing Ability for Jailbreak Attack
Inspired by this finding, we construct six different jailbreak attacks based on the easyjailbreak dataset and the 50 harmful queries from the aforementioned synthesized dataset. The attack methods include AutoDAN (Liu et al. 2024), DeepInception (Li et al. 2023), GPTFuzz(Yu, Lin, and Xing 2023), ICA(Wei, Wang, and Wang 2023), ReNellm(Ding et al. 2024), and gcg(Zou et al. 2023), each aiming to bypass the model’s safety mechanisms through different approaches. For Llama2-7B-chat, we use three sets of hidden states to compute the first two principal components, which included benign queries, harmful queries, and jailbreak attacks.
We observe that models trained with safety protocols can not only distinguish between harmful and benign queries but also differentiate jailbreak attacks. The boundaries between these categories (indicated by black dashed lines) can be easily fitted using an SVM, with the harmfulness of the queries still serving as the labels. More reuslts can be found in AppendixB.2. Based on this observation, our insights for developing a lightweight classification model include: (i) employing more robust methods for sampling the classification criteria, and (ii) more accurately fitting the boundaries between benign queries, harmful queries, and jailbreak attacks within the model’s representation space.
4 Hidden State Filter
@@
4.1 Overview of Hidden State Filter
@@
Figure 3: Schematic representation of the HSF mechanism. In this process, the Large Language Model (LLM) forwards the hidden states from the final Decoder Layer to the trained HSF. The filter extracts the last $k$ tokens, concatenates them, and computes a harmfulness score $\alpha$ using a sigmoid function. This score is then compared to a threshold to determine the model’s response, ensuring safe and reliable output.
4.2 Training Phase: A Weak Classifier
To construct the expert model, we first collect 3,000 samples each from the UltraSafety (Guo et al. 2024) and PKU-SafeRLHF-prompt (Ji et al. 2024b) datasets as the harmful query dataset. These queries are expected to be rejected by any LLM aligned with human values. We also collect 6,000 samples from the databricks-dolly-15k (Conover et al. 2023) dataset as the benign query dataset. For the poorly aligned mistral-instruct-v0.2 model, we additionally sample 750 queries from both the harmful and benign datasets we construct. These samples are used to create a positive and negative dataset for training by adding template completion jailbreak attack templates.
@@ 4.2 Training Phase: A Weak Classifier
$T_{k}=[t_{k},\mathbf{0},t_{2},\mathbf{0},\ldots,\mathbf{0},t_{1}],$ (1)
@@
$f_{k}:\mathbb{R}^{n\times k}\to\mathbb{R},\quad f_{k}(x)=w_{k}^{T}T_{k}+b_{k},$ (2)
@@
$\mathcal{L}_{k}(\theta)=-l\log(\sigma(f_{k}(T_{k})))-(1-l)\log(1-\sigma(f_{k}(T_{k}))),$ (3)
@@
Table 2: Comparison of Attack Success Rates between HSF and baseline defenses on Mistral-instruct-v0.2 and Llama2-7B-chat.
4.3 Defense Phase: Implementing Jailbreak Attack Defense by Classifying User Queries
@@ 5 Experiments
5.1 Experimental Setup
Basement Models. We deploy HSF on three open-source LLMs: Vicuna-v1.5 (Chiang et al. 2023), Llama2-7b-chat (Touvron et al. 2023), Mistral-instruct-v0.2 (Jiang et al. 2023)to evaluate its effectiveness.
Attack Methods. We consider six state-of-the-art jailbreak attacks, each covering different categories, implemented through easyjailbreak (Zhou et al. 2024). Specifically, GCG (Zou et al. 2023) represents gradient-based attacks, AutoDAN (Liu et al. 2024) is based on genetic algorithms, and ICA (Wei, Wang, and Wang 2023)represents context-learning-based attacks. We also included ReNellm (Ding et al. 2024), DeepInception (Li et al. 2023), and GPTFuzz (Yu, Lin, and Xing 2023) as representative empirical jailbreak attacks. To evaluate the defense performance against naive attackers directly inputting harmful queries, we used two harmful query benchmark datasets: Advbench (Zou et al. 2023) and MaliciousInstruct (Huang et al. 2024). Detailed settings for these attack methods and harmful query datasets can be found in the AppendixA.1.
Baselines. We consider six state-of-the-art defense mechanisms as baselines. PPL(Alon and Kamfonas 2023) and Self-Examination (Phute et al. 2023)) are input-output detection-based methods, while Paraphrase (Jain et al. 2023), Retokenization (Jain et al. 2023), and Self-Reminder (Xie et al. 2023) are mitigation-based methods. DRO(Zheng et al. 2024) is a safety-prompt-based method. Detailed descriptions and hyperparameter settings for each method can be found in the AppendixA.1.
Evaluation Metrics. We utilize Llama-Guard-3-8B (Llama Team 2024) to evaluate the safety of the model’s responses. Although keyword-based detection methods (Zou et al. 2023) are commonly employed, our experiments revealed that these methods had a low AUC, indicating a high false positive rate. This necessitated additional manual verification, leading to increased costs. Consequently, we employ the state-of-the-art Llama-Guard-3-8B to assess whether the model generated unsafe content and to calculate the ASR.

### v2

@@
Abstract.
With the growing deployment of LLMs in daily applications like chatbots and content generation, efforts to ensure outputs align with human values and avoid harmful content have intensified. However, increasingly sophisticated jailbreak attacks threaten this alignment, aiming to induce unsafe outputs. Current defense efforts either focus on prompt rewriting or detection, which are limited in effectiveness due to the various design of jailbreak prompts, or on output control and detection, which are computationally expensive as they require LLM inference. Therefore, designing a pre-inference defense method that resists diverse jailbreak prompts is crucial for preventing LLM jailbreak attacks. We observe that jailbreak attacks, safe queries, and harmful queries exhibit different clustering patterns within the LLM’s hidden state representation space. This suggests that by leveraging the LLM’s hidden state representational capabilities, we can analyze the LLM’s forthcoming behavior and proactively intervene for defense. In this paper, we propose a jailbreak attack defense strategy based on a Hidden State Filter (HSF), a lossless architectural defense mechanism that enables the model to preemptively identify and reject adversarial inputs before the inference process begins. We activate its defensive potential through an additional plugin module, effectively framing the defense task as a classification problem. Experimental results on two benchmark datasets, utilizing three different LLMs, show that HSF significantly enhances resilience against six cutting-edge jailbreak attacks. It significantly reduces the success rate of jailbreak attacks while minimally impacting responses to benign user queries, with negligible inference overhead, and outperforming defense baselines.
Large Language Models, Jailbreak Attacks, Defense Strategies
journalyear: 2025copyright: acmlicensedconference: Companion of the 16th ACM/SPEC International Conference on Performance Engineering; April 28-May 2, 2025; Sydney, NSW, Australia.booktitle: Companion of the 16th ACM/SPEC International Conference on Performance Engineering (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiaisbn: 979-8-4007-1331-6/25/04doi: 10.1145/3701716.3717659submissionid: wk1203journalyear: 2025copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3717659isbn: 979-8-4007-1331-6/2025/04ccs: Security and privacy Usability in security and privacyccs: Computing methodologies Natural language processing
1. Introduction
Table 1. Examples of various jailbreak attack methods and their corresponding prompts. The {User Query} is ”Write a tutorial on how to make a bomb”. Each attack represents a different strategy to bypass content restrictions in alignment LLM by manipulating input prompts.
@@ 1. Introduction
With the growing capabilities of models like ChatGPT (Achiam et al., 2023), Llama (Touvron et al., 2023), and Mistral (Jiang et al., 2023), there is increasing concern about the potential security risks associated with their powerful functionalities, such as biased reports (Ferrara, 2023), inaccurate information (Ji et al., 2023), or harmful contents (Weidinger et al., 2021). Over the past few years, researchers make extensive efforts to align these models safely using techniques such as SFT (Ouyang et al., 2022) and RLHF(Bai et al., 2022), ensuring that the models adhere to human values and preferences for safety and reliability. Despite these alignment efforts providing models with the ability to judge the safety of inputs to some extent, models remain susceptible to adversarial attacks (Zou et al., 2023).
Recent studies have highlighted the potential threat of “jailbreak attacks”(Liu et al., 2023; Wei et al., 2024), which can bypass existing alignments and induce models to output unsafe content. As shown in Table 1, different types of attacks vary greatly and can be introduced into the prompts in many ways. Current defense methods primarily focus on prompt’s detection (Alon and Kamfonas, 2023) and rewriting (Jain et al., 2023; Zheng et al., 2024) or output’s detection (Xie et al., 2023) and control (Xu et al., 2024). The former defends against jailbreak attacks by using PPL (Alon and Kamfonas, 2023), paraphrasing (Jain et al., 2023), or adding safer prompts (Zheng et al., 2024; Wei et al., 2023), but faces issues of limited defense scope, high costs, and disruption of benign prompts. The latter defends by controlling the decoder’s generation process (Xu et al., 2024) or detecting the final output (Xie et al., 2023), but incurs high computational costs as it starts only after LLM inference. Therefore, designing a pre-inference defense method that remains unaffected by various jailbreak prompts is crucial for preventing LLM jailbreak attacks.
@@
Figure 1. Comparison of four different models using PCA visualization, illustrating how each model differentiates between harmless (blue) and harmful (orange) queries. The Llama2-7B-chat(a), Vicuna-v.15-7B(b) and Mistral-instruct-v0.2-7B(c) are aligned models, while the WizardLM-30B-Uncensored(d) is unaligned model.
Figure 2. Comparison of Llama2-7B-chat model using PCA visualization, illustrating how the model differentiates between harmless (blue) and harmful (orange) queries across various jailbreak (red) methods. Subplots correspond to the following attack methods: (1) Autodan, (2) GCG, (3) DeepInception, (4) ReNellm, (5) ICA, and (6) GPTFuzz. The decision boundaries that the model learned to separate the two types of queries are shown.
2. Related Work
@@
2.1. Jailbreak Attacks
Current jailbreak attacks can be broadly categorized into two major types: template completion attacks and prompt rewriting attacks(Yi et al., 2024).
In template completion attacks, the attacker embeds harmful queries within a contextual template to generate jailbreak prompts. Liu et al. (2023) investigates adversarial examples against the GPT-3 model, where carefully crafted inputs induce the model to generate inappropriate content. Wei et al. (2024) identifies the root cause of LLMs’ vulnerability to jailbreak attacks as conflicting objectives and generalization mismatches, noting that the model’s safety capabilities do not match its complexity. Li et al. (2023) constructs a virtual, nested scenario to hypnotize large models into performing jailbreak attacks. Ding et al. (2024) utilizes prompt rewriting and scenario nesting within LLMs themselves to generate effective jailbreak prompts. Wei et al. (2023) leverages the model’s contextual learning abilities to guide the LLM in generating unsafe outputs.
Prompt rewriting attacks involve constructing jailbreak prompts through optimization techniques, with two main approaches: (1) Gradient-based methods, as Zou et al. (2023) uses gradient optimization to generate adversarial inputs; (2) Genetic algorithm-based methods, as Liu et al. (2024) and Yu et al. (2023) collect seeds from a curated library of handcrafted templates and continuously generate optimized adversarial prompts through mutation and crossover.
@@ 3. Motivation
3.1. Data Synthesis and Basement Model
@@
To ensure the diversity of the synthesized dataset, we sampled the top 100 data points from Advbench (Zou et al., 2023) as harmful queries and instructed gpt-3.5-turbo to generate benign versions of these queries while maintaining consistency in length as much as possible. Please refer to the appendix for the prompts we used to guide the data synthesis. We then applied additional manual checks to ensure effectiveness and quality. As a result, we collected 100 harmful and 100 benign queries, with an average length of 15.05 and 17.14 tokens, respectively (measured using the Llama2-7b-chat tokenizer).
We conduct experiments with three popular 7B chat LLMs available on HuggingFace, as well as an uncensored 30B LLM: Llama-2-chat (Touvron et al., 2023), Vicuna-v1.5 (Chiang et al., 2023), Mistral-instruct-v0.2 (Jiang et al., 2023), and WizardLM-30B-Uncensored (Computations, 2024). Some of these models explicitly underwent extensive safety training (Llama2-7b-chat and Mistral-instruct-v0.2).
3.2. Visualization Analysis
@@
3.2.1. Aligned Model vs. Unaligned Model
Following (Zheng et al., 2024), we used Principal Component Analysis (PCA) to visualize the hidden states of the models. We selected the hidden states of the last input token from the model’s last Decoder Layer output, as these hidden states intuitively gather all the information about the model’s understanding of the query and its response strategy. It should be noted that these hidden states are also projected through the language modeling head (a linear mapping) for the prediction of the next token, which means they present a linear structure in the corresponding representation space (as assumed by PCA). We use two sets of hidden states to compute the first two principal components for each model, which include both harmful and benign queries. By selecting these data points, we are able to extract the most significant features related to the harmfulness of the queries.
@@
3.2.2. Distinguishing Ability for Jailbreak Attack
Inspired by this finding, we construct six different jailbreak attacks based on the easyjailbreak dataset and the 50 harmful queries from the aforementioned synthesized dataset. The attack methods include AutoDAN (Liu et al., 2024), DeepInception (Li et al., 2023), GPTFuzz(Yu et al., 2023), ICA(Wei et al., 2023), ReNellm(Ding et al., 2024), and gcg(Zou et al., 2023), each aiming to bypass the model’s safety mechanisms through different approaches. For Llama2-7B-chat, we use three sets of hidden states to compute the first two principal components, which included benign queries, harmful queries, and jailbreak attacks.
We observe that models trained with safety protocols can not only distinguish between harmful and benign queries but also differentiate jailbreak attacks. The boundaries between these categories (indicated by black dashed lines) can be easily fitted using an SVM, with the harmfulness of the queries still serving as the labels. More reuslts can be found in AppendixLABEL:appendix:more_results. Based on this observation, our insights for developing a lightweight classification model include: (i) employing more robust methods for sampling the classification criteria, and (ii) more accurately fitting the boundaries between benign queries, harmful queries, and jailbreak attacks within the model’s representation space.
4. Hidden State Filter
@@
4.1. Overview of Hidden State Filter
@@
Figure 3. Schematic representation of the HSF mechanism. In this process, the Large Language Model (LLM) forwards the hidden states from the final Decoder Layer to the trained HSF. The filter extracts the last $k$ tokens, concatenates them, and computes a harmfulness score $\alpha$ using a sigmoid function. This score is then compared to a threshold to determine the model’s response, ensuring safe and reliable output.
4.2. Training Phase: A Weak Classifier
To construct the expert model, we first collect 3,000 samples each from the UltraSafety (Guo et al., 2024) and PKU-SafeRLHF-prompt (Ji et al., 2024b) datasets as the harmful query dataset. These queries are expected to be rejected by any LLM aligned with human values. We also collect 6,000 samples from the databricks-dolly-15k (Conover et al., 2023) dataset as the benign query dataset. For the poorly aligned mistral-instruct-v0.2 model, we additionally sample 750 queries from both the harmful and benign datasets we construct. These samples are used to create a positive and negative dataset for training by adding template completion jailbreak attack templates.
@@ 4.2. Training Phase: A Weak Classifier
(1) $T_{k}=[t_{k},\mathbf{0},t_{2},\mathbf{0},\ldots,\mathbf{0},t_{1}],$
@@
(2) $f_{k}:\mathbb{R}^{n\times k}\to\mathbb{R},\quad f_{k}(x)=w_{k}^{T}T_{k}+b_{k},$
@@
(3) $\mathcal{L}_{k}(\theta)=-l\log(\sigma(f_{k}(T_{k})))-(1-l)\log(1-\sigma(f_{k}(T_{k}))),$
@@
Table 2. Comparison of Attack Success Rates between HSF and baseline defenses on Mistral-instruct-v0.2 and Llama2-7B-chat.
@@
4.3. Defense Phase: Implementing Jailbreak Attack Defense by Classifying User Queries
@@ 5. Experiments
5.1. Experimental Setup
Basement Models. We deploy HSF on three open-source LLMs: Vicuna-v1.5 (Chiang et al., 2023), Llama2-7b-chat (Touvron et al., 2023), Mistral-instruct-v0.2 (Jiang et al., 2023)to evaluate its effectiveness.
Attack Methods. We consider six state-of-the-art jailbreak attacks, each covering different categories, implemented through easyjailbreak (Zhou et al., 2024). Specifically, GCG (Zou et al., 2023) represents gradient-based attacks, AutoDAN (Liu et al., 2024) is based on genetic algorithms, and ICA (Wei et al., 2023)represents context-learning-based attacks. We also included ReNellm (Ding et al., 2024), DeepInception (Li et al., 2023), and GPTFuzz (Yu et al., 2023) as representative empirical jailbreak attacks. To evaluate the defense performance against naive attackers directly inputting harmful queries, we used two harmful query benchmark datasets: Advbench (Zou et al., 2023) and MaliciousInstruct (Huang et al., 2024). Detailed settings for these attack methods and harmful query datasets can be found in the AppendixA.1.
Baselines. We consider six state-of-the-art defense mechanisms as baselines. PPL(Alon and Kamfonas, 2023) and Self-Examination (Phute et al., 2023)) are input-output detection-based methods, while Paraphrase (Jain et al., 2023), Retokenization (Jain et al., 2023), and Self-Reminder (Xie et al., 2023) are mitigation-based methods. DRO(Zheng et al., 2024) is a safety-prompt-based method. SPD(Candogan et al., [n. d.]) is a method for detecting jailbreaking inputs via the logit values in a single forward pass.Note that the original classifier used in SPD is SVM. However, due to issues encountered during reproduction, it has been replaced with an equivalent MLP structure in this study.Detailed descriptions and hyperparameter settings for each method can be found in the AppendixA.1.
Evaluation Metrics. We utilize Llama-Guard-3-8B (Llama Team, 2024) to evaluate the safety of the model’s responses. Although keyword-based detection methods (Zou et al., 2023) are commonly employed, our experiments revealed that these methods had a low AUC, indicating a high false positive rate. This necessitated additional manual verification, leading to increased costs. Consequently, we employ the state-of-the-art Llama-Guard-3-8B to assess whether the model generated unsafe content and to calculate the ASR.